# A Practical Adaptive Subgame Perfect Gradient Method

### v1

@@ 3.5 Preconditioning
$B_{i+1}=(I-\frac{s_{i}y_{i}^{T}}{y_{i}^{T}s_{i}})B_{i}(I-\frac{s_{i}y_{i}^{T}}{y_{i}^{T}s_{i}})+\frac{s_{i}s_{i}^{T}}{y_{i}^{T}s_{i}},\qquad\qquad B_{i+1}^{-1}=B_{i}^{-1}-\frac{B_{i}^{-1}s_{i}s_{i}^{T}B_{i}^{-1}}{s_{i}^{T}B_{i}^{-1}s_{i}}+\frac{y_{i}y_{i}^{T}}{y_{i}^{T}s_{i}}$ (3.19)
@@
 $y={x}^{(\ell)}_{0}+10^{-4}\cdot\mathcal{N}(0,1)$
@@
In total, ASPGM requires the storage of $3k+2t$ additional vectors of size $d$ (beyond the current $x_{n}$ and $g_{n}$), where $k$ is the memory size used in the BSPGM subalgorithm, and $t$ is the memory size used for preconditioning. All other storage requirements are negligible at high dimensions. In comparison, the storage requirements of SPGM and L-BFGS are $2k$ and $2t$, respectively.
@@ 4.1 Algorithm Varieties and Benchmarks
At the end of our section, we compare in Figure˜6 the performance of ASPGM with different memory sizes. As expected, we see diminishing returns as the memory size increases. This is consistent with many memory-based methods including SPGM and L-BFGS. These results justify our use of ASPGM-5-5 as our default method. Even at small memory, we see performance is generally maintained, motivating our inclusion of ASPGM-1-1 in our results as a very low-memory alternative.
@@
We compare the performance of our base BSPGM method and our full ASPGM method with a wide range of competing algorithms. We focus primarily on adaptive methods and thus exclude methods requiring smoothness estimates, like OGM and SPGM, from our survey. Many of these methods require an unspecified initial estimate of $L_{0}$; we initialize $L_{0}=\tilde{L}_{I}(x_{0},x_{0}+10^{-4}\cdot\xi)$ for $\xi\in\mathcal{N}(0,1)^{d}$. L-BFGS is implemented in Julia using the Optim [46] package. All other algorithms are implemented directly in Julia.
@@
  • AdaNAG: AdaNAG-G12 from [8, Corollary 7].
@@
  • AdGD: Adaptive Gradient Descent [10]. We use the improved adaptive gradient descent method from [10, Algorithm 2].
  • NAGF: NAG-free algorithm from [12, Algorithm 2].
  • OSGMR: Online Scaled Gradient Method with ratio surrogate [19], with AdaGrad as the preconditioner update algorithm. For the stepsize, we conducted a parameter sweep across a subset of problems and selected $\eta=0.1$ as the value with the best overall performance. Improved performance would follow from tuning this parameter per problem instance. We follow the authors’ approach for handling an unknown lower bound (See [19, Section B.1]).
@@ 4.2 Synthetic Smooth Convex Problems
We generate random problem instances for several problem classes discussed below, all parameterized by $A\in\mathbb{R}^{m\times d}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{m}$, and $x_{0}\in\mathbb{R}^{d}$. For each problem instance we set $m=4d$, $x_{0}=0$, generate $b$ element-wise from $\mathcal{N}(0,1)$, and generate $c$ element-wise from $\mathcal{U}\{0,1\}$. Since the conditioning of each problem class is determined by the condition number of $A^{T}A$, we construct $A$ so that $A^{T}A$ has approximate condition numbers of $\kappa=10^{2}$ and $\kappa=10^{4}$. We also distribute the singular values of $A$ in two different ways:
@@
  • Bimodal distribution: $\sigma_{1},\dots,\sigma_{9d/10}\sim\mathcal{U}(1,1.1)$ and $\sigma_{9d/10+1},\dots,\sigma_{d}\in\mathcal{U}(0.9\sqrt{\kappa},\sqrt{\kappa}).$
We generate 4 instances for each condition number and each distribution above; using dimensions $d=1000,2000,4000,8000$ and 6 problem classes below, this yields 384 synthetic problem instances.
@@
In Figure˜1, we present the overall performance of ASPGM, BSPGM, and competing methods across all of our problem classes. We also include in Section˜A.3 these same performance results stratified by problem class. We see ASPGM with both memory size $5$ and a minimal size of $1$ is highly performant on our synthetic problem set, outperforming other adaptive methods, as well as LBFGS-HZ. This advantage is maintained when considering wall clock time. We also note the strong performance of BSPGM for attaining low-accuracy solutions, especially in terms of wall-clock time. This is particularly impressive since BSPGM also provides an explicit non-asymptotic convergence guarantee and corresponding stopping criteria via $L_{n}/\tau_{n}$.
@@ 4.3 Real-Data Regression Problems
We also present performance results for problem instances on real data. Using the LIBSVM [47] library, we apply least squares regression (4.1) to the data sets bodyfat, eunite2001, pyrim, triazines, and yearPredictionMSD and logistic regression with L2 regularization (4.2) to the data sets colon-cancer, duke-breast-cancer, gisette, leukemia, and madelon. Results are shown in Figure˜2. We see similar trends as in our synthetic problems, but see that ASPGM no longer outperforms LBFGS-HZ. We also notice a stronger difference for ASPGM between oracle performance and real time performance. This can likely be partially attributed to these problems having lower dimension, thereby causing our subproblem to have a larger proportional impact on computation time. In Figure˜3, we show detailed results for the colon-cancer logistic regression problem. Here we show the modified performance measure $(f(x_{n})-f_{\star})/(\frac{1}{2}\|x_{0}-x_{\star}\|^{2})$, so we can directly compare the performance with the guarantees (ignoring error terms) of BSPGM-7, OBL, and AdaNAG in the center plot. We also show the value of the error term $\Delta_{n}$ for BSPGM-7 and OBL. In particular, this illustrates our dynamic theory allowing $\Delta_{n}$ to briefly fluctuate and then returns to nearly zero for BSPGM-7, unlike OBL.
@@
Figure 3: Performance data from logistic regression (4.2) run on the colon-cancer dataset. On the left, we show the alternate performance measure $(f(x_{n})-f_{\star})/(\frac{1}{2}\|x_{0}-x_{\star}\|^{2})$. In the middle plot, we compare the guarantees of BSPGM-7, OBL, and AdaNAG (ignoring error terms). For BSPGM-7 and OBL, this corresponds to $L_{n}/\tau_{n}$. The right plot shows the accumulated error $\Delta_{n}$ for BSPGM-7 and OBL.
@@ 4.4 Real-Data Feasibility Problems
As another set of problems derived from real data, we consider the search for a feasible point of an LP. Given an LP with the constraints $Ax\leq b$, one can frame this objective as minimizing $\max_{i}(Ax-b)_{i}$. If we apply either of our smoothing approaches (4.3), (4.4) from Section˜4.2, we obtain a smooth problem suitable for our set of algorithms. We apply this method to the following problems from the “LPfeas Benchmark” dataset [48]: brazil3, chromaticindex1024-7, ex10, graph40-40, qap15, savsched1, scpm1, set-cover-model, and supportcase10. These include high-dimensional ($d\approx 10^{5}$) sparse datasets. We apply both smoothing approaches to each data file, obtaining 18 total problem instances. We see similar trends in oracle-call performance but diminished real-time performance. We attribute this to the sparsity of data matrices, making gradient calculations more efficient and increasing the relative cost of our subproblem calculations.
@@ 4.5 Poorly Conditioned Problems
Lastly, to compare performances in highly degenerate settings, we consider several known poorly conditioned problems from the literature. Each problem is a quadratic of the form $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$, with specially selected $A\in\mathbb{R}^{d\times d}$, $b\in\mathbb{R}^{d}$, and starting point $x_{0}\in\mathbb{R}^{d}$. First, we consider the classic hard instance [49] with $x_{0}=0$, $b=(-\frac{1}{2},0,\dots,0)$ and $A$ as the tridiagonal matrix with diagonal $(1,\dots,1)$ and super- and sub-diagonals $(-\frac{1}{2},\dots,-\frac{1}{2})$. Next, we consider from [25], the case where $b=0$, AA is diagonal entries with entries $a_{ii}=\sin^{2}(\frac{\pi i}{2d})$, and $x_{0}=(1/a_{11},\dots,1/a_{dd})$. Finally, we consider the instance with $x_{0}=0$, $b=(1,\dots,1)$ and $A$ as the diagonal matrix with diagonal $(1,2,\dots,d)$. We refer to these problem instances as Problem A, Problem B, and Problem C, respectively. In Figure˜5, we show our methods remain competitive in these degenerate settings with $d=1000$.
@@ 5 Theory
$\epsilon(ce_{s},0)$ $=ca_{s}+\delta_{n}-\frac{L_{n}}{2}\|z_{s+1}-x_{0}\|^{2}=c\tau_{s}(f_{s}-\frac{1}{2L_{s}}\|g_{s}\|^{2}-v_{m})+\delta_{n}=c\tau_{s}(v_{s}-v_{m})+\delta_{n}\geq 0$
@@
$=\kappa\tau_{s}f_{s}-\kappa\tau_{s}\frac{1}{2L_{s}}\|{g_{s}}\|^{2}-\frac{\kappa L_{s}}{2}\|{x_{0}}\|^{2}+\frac{\kappa L_{s}}{2}\|{z_{s+1}}\|^{2}-\kappa v_{m}\tau_{s}-L_{n}c\langle{z_{s+1}-x_{0}},{x_{0}}\rangle$
$\qquad\qquad\qquad\qquad+L_{n}\tau_{s}\left(\frac{1}{L_{s}^{2}}-\frac{1}{L_{n}^{2}}\right)\frac{1}{2}\|{g_{s}}\|^{2}-\frac{L_{n}}{2}\|{c(z_{s+1}-x_{0})}\|^{2}$
$\geq\kappa\tau_{s}f_{s}-\kappa\tau_{s}\frac{1}{2L_{n}}\|{g_{s}}\|^{2}-\kappa\tau_{s}v_{m}+\tau_{s}\left(\frac{1}{L_{s}}-\frac{1}{L_{n}}\right)\frac{1}{2}\|g_{s}\|^{2}+\frac{L_{n}(c-c^{2})}{2}\|z_{s+1}-x_{0}\|^{2}$
@@
$U^{\prime}$ $=\left(\frac{L_{n}}{2}\|{x_{0}}\|^{2}-\frac{L_{n}}{2}\|{z^{\prime}}\|^{2}-\tau^{\prime}v_{m}+\Delta^{\prime}+\delta_{n}\right)+\tau^{\prime}f_{\star}+\langle L_{n}(z^{\prime}-x_{0}),x_{\star}\rangle,$ (3.7)
$U^{\prime}$ $=\left(-\sum_{i={n-k}}^{n-1}{\rho}_{i}{h}_{i}-\sum_{i={n-k}}^{n-1}{\gamma}_{i}{w}_{i}+\sum_{i={n-k}}^{n-1}{\rho}_{i}{\Delta}_{i}+\epsilon\right)+\left(\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}\right)f_{\star}+\langle L_{n}Z\rho-L_{n}G\gamma,x_{\star}\rangle.$ (3.8)
@@
$\left\{\begin{array}[]{ll}\max_{\rho,\gamma}&\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}\\ \text{s.t.}&\frac{L_{n}}{2}\|{Z\rho-G\gamma}\|^{2}\leq\sum_{i=n-k}^{n-1}\rho_{i}(h_{i}-v_{m}\tau_{i})+\sum_{i=n-k}^{n-1}\gamma_{i}(w_{i}-v_{m})-\langle L_{n}Z\rho,x_{0}\rangle+\langle L_{n}G\gamma,x_{0}\rangle+\delta_{n}\\ &\rho,\,\gamma\geq 0.\end{array}\right$. (5.20)
Recalling our definition of $\epsilon(\rho,\gamma)$ from (3.11), this is exactly our optimization problem (3.12). Finally, with $\tau^{\prime}$ as the solution to (5.20), i.e. $\tau^{\prime}=\lambda_{n}=\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}$, and $\tau_{n}$ set according to Algorithm˜3, we recover the exact form of Algorithm˜4.
@@
The final equality of our theorem $\frac{1}{\tau_{0,N}}=\frac{2}{N(N+1)+\sqrt{2N(N+1)}}$ comes directly from [38, Theorem 4].
@@
[…]
$\geq\frac{2{L}^{(\ell)}_{n+1}}{\mu}+\frac{{L}^{(\ell)}_{n+1}{\Delta}^{(\ell)}_{n+1}}{f({x}^{(\ell)}_{0})-f(x_{\star})}$ (5.22)
$=\frac{2{L}^{(\ell)}_{n+1}(f({x}^{(\ell)}_{0})-f(x_{\star}))+\mu{L}^{(\ell)}_{n+1}{\Delta}^{(\ell)}_{n+1}}{\mu(f({x}^{(\ell)}_{0})-f(x_{\star}))}.$ (5.23)
@@
$\leq\frac{{L}^{(\ell)}_{n+1}}{{\tau}^{(\ell)}_{n+1}}\left(\frac{f({x}^{(\ell)}_{0})-f(x_{\star})}{\mu}+\frac{{\Delta}^{(\ell)}_{n+1}}{2}\right)$

### v2

@@ 3.5 Preconditioning
$B_{i+1}=(I-\frac{s_{i}y_{i}^{T}}{y_{i}^{T}s_{i}})B_{i}(I-\frac{y_{i}s_{i}^{T}}{y_{i}^{T}s_{i}})+\frac{s_{i}s_{i}^{T}}{y_{i}^{T}s_{i}},\qquad\qquad B_{i+1}^{-1}=B_{i}^{-1}-\frac{B_{i}^{-1}s_{i}s_{i}^{T}B_{i}^{-1}}{s_{i}^{T}B_{i}^{-1}s_{i}}+\frac{y_{i}y_{i}^{T}}{y_{i}^{T}s_{i}}$ (3.19)
@@
 $y={x}^{(\ell)}_{0}-10^{-4}\cdot\frac{\nabla f({x}^{(\ell)}_{0})}{\|\nabla f({x}^{(\ell)}_{0})\|}$
@@
In total, ASPGM requires the storage of $3k+2t$ additional vectors of size $d$ (beyond $x_{0}$, $x_{n}$, and $g_{n}$), where $k$ is the memory size used in the BSPGM subalgorithm, and $t$ is the memory size used for preconditioning. All other storage requirements are negligible at high dimensions. In comparison, the storage requirements of SPGM and L-BFGS are $2k$ and $2t$, respectively.
@@ 4.1 Algorithm Varieties and Benchmarks
At the end of our section, we compare in Figure˜6 the performance of ASPGM with different memory sizes. As expected, we see diminishing returns as the memory size increases. This is consistent with many memory-based methods including SPGM and L-BFGS. These results justify our use of ASPGM-5-5 as our default method. At memory size 1, the subproblem (3.12) can be solved explicitly, eliminating the need for an SOCP solve. This enabled us to provide a stand-alone implementation of ASPGM-1-1 in our repository above, with no dependencies on external solvers. Our results show that even with very low memory, ASPGM-1-1 generally maintains performance, with wall-clock performance notably benefiting from the highly efficient explicit solve.
@@
We compare the performance of our base BSPGM method and our full ASPGM method with a wide range of competing algorithms. We focus primarily on adaptive methods and thus exclude methods requiring smoothness estimates, like OGM and SPGM, from our survey. Many of these methods require an unspecified initial estimate of $L_{0}$; we initialize $L_{0}=\tilde{L}_{I}(x_{0},x_{0}-10^{-4}\cdot\frac{\nabla f(x_{0})}{\|\nabla f(x_{0})\|})$. L-BFGS is implemented in Julia using the Optim [36] package. All other algorithms are implemented directly in Julia.
@@
  • AdaNAG: AdaNAG-G12 from [50, Corollary 7].
@@
  • AdGD: Adaptive Gradient Descent [34]. We use the improved adaptive gradient descent method from [34, Algorithm 2].
  • NAGF: NAG-free algorithm from [8, Algorithm 2].
  • OSGMB: OSGM-Best method from [10], a variant of Online Scaled Gradient Method [17]. We follow the implementation of the authors and set parameters $\omega=0$, $\tau=\frac{1}{2}L_{0}^{2}$, $\beta_{0}=0.95$, $\eta_{P}=\frac{1}{L_{0}}$, and $\eta_{\beta}=\min(1.0,L_{0})$.
@@
We employ an oracle model in which function and gradient information are returned simultaneously (i.e., $x\mapsto(f(x),\nabla f(x))$). Each such evaluation is considered an oracle call in our plots below. For algorithms that consider function oracles and gradient oracles separately, we apply special handling to efficiently leverage the joint first-order oracle.
@@ 4.2 Synthetic Smooth Convex Problems
We generate random problem instances for several problem classes discussed below, all parameterized by $A\in\mathbb{R}^{p\times d}$, $b\in\mathbb{R}^{p}$, $c\in\mathbb{R}^{p}$, and $x_{0}\in\mathbb{R}^{d}$. For each problem instance we set $p=4d$, $x_{0}=0$, generate $b$ element-wise from $\mathcal{N}(0,1)$, and generate $c$ element-wise from $\mathcal{U}\{0,1\}$. Since the conditioning of each problem class is determined by the condition number of $A^{T}A$, we construct $A$ so that $A^{T}A$ has approximate condition numbers of $\kappa=10^{2}$ and $\kappa=10^{4}$. We also distribute the singular values of $A$ in two different ways:
@@
  • Bimodal distribution: $\sigma_{1},\dots,\sigma_{9d/10}\sim\mathcal{U}(1,1.1)$ and $\sigma_{9d/10+1},\dots,\sigma_{d}\sim\mathcal{U}(0.9\sqrt{\kappa},\sqrt{\kappa}).$
We generate two instances for each condition number and each distribution above; using dimensions $d=1000,2000,4000,8000$ and 6 problem classes below, this yields 192 synthetic problem instances.
@@
In Figure˜1, we present the overall performance of ASPGM, BSPGM, and competing methods across all of our problem classes. We also include in Section˜A.3 these same performance results stratified by problem class. We see ASPGM with both memory size $5$ and a minimal size of $1$ is highly performant on our synthetic problem set, outperforming other adaptive methods, as well as LBFGS-HZ. This advantage is maintained when considering wall clock time. We also note the strong performance of BSPGM for attaining low-accuracy solutions, especially in terms of wall-clock time. This is particularly impressive since BSPGM also provides an explicit non-asymptotic convergence guarantee and corresponding stopping criterion via $L_{n}/\tau_{n}$.
@@ 4.3 Real-Data Regression Problems
We also present performance results for problem instances on real data. Using the LIBSVM [9] library, we apply least squares regression (4.1) to the data sets bodyfat, eunite2001, pyrim, triazines, and yearPredictionMSD and logistic regression with L2 regularization (4.2) to the data sets colon-cancer, duke-breast-cancer, gisette, leukemia, and madelon. Results are shown in Figure˜2; for clarity we include only a subset of our methods, focusing on those with the strongest performance throughout our experiments. We see similar trends as in our synthetic problems, but see that ASPGM no longer outperforms LBFGS-HZ. We also notice a stronger difference for ASPGM between oracle performance and wall-clock performance. This can be partially attributed to these problems having lower dimension, thereby causing our subproblem to have a larger proportional impact on computation time. In Figure˜3, we show detailed results for the duke-breast-cancer logistic regression problem. Here we show the modified performance measure $(f(x_{n})-f_{\star})/(\frac{1}{2}\|x_{0}-x_{\star}\|^{2})$, so we can directly compare the performance with the guarantees (ignoring error terms) of BSPGM-7 and OBL in the center plot. We also show the value of the error term $\Delta_{n}$ for BSPGM-7 and OBL. In particular, this illustrates our dynamic theory allowing $\Delta_{n}$ to briefly fluctuate and then return to nearly zero for BSPGM-7, unlike OBL.
@@
Figure 3: Performance data from logistic regression (4.2) run on the colon-cancer dataset. On the left, we show the alternate performance measure $(f(x_{n})-f_{\star})/(\frac{1}{2}\|x_{0}-x_{\star}\|^{2})$. In the middle plot, we compare the guarantees of BSPGM-7 and OBL (ignoring error terms). This corresponds to $L_{n}/\tau_{n}$. The right plot shows the accumulated error $\Delta_{n}$ for BSPGM-7 and OBL.
@@ 4.4 Real-Data Feasibility Problems
As another set of problems derived from real data, we consider the search for a feasible point of an LP. Given an LP with the constraints $Ax\leq b$, one can frame this objective as minimizing $\max_{i}\{(Ax-b)_{i},0\}$. If we apply either of our smoothing approaches (4.3) or (4.4) from Section˜4.2, we obtain a smooth problem suitable for our set of algorithms. We apply this method to the following problems from the “LPfeas Benchmark” dataset [35]: brazil3, chromaticindex1024-7, ex10, graph40-40, qap15, savsched1, scpm1, set-cover-model, and supportcase10. These include high-dimensional ($d\approx 10^{5}$) sparse datasets. We apply both smoothing approaches to each data file, obtaining 18 total problem instances. We see similar trends in oracle-call performance but reasonable gains in wall-clock performance.
@@ 4.5 Poorly Conditioned Problems
Lastly, to compare performance in highly degenerate settings, we consider several known poorly conditioned problems from the literature. Each problem is a quadratic of the form $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$, with specially selected $A\in\mathbb{R}^{d\times d}$, $b\in\mathbb{R}^{d}$, and starting point $x_{0}\in\mathbb{R}^{d}$. First, we consider the classic hard instance [41] with $x_{0}=0$, $b=(-\frac{1}{2},0,\dots,0)$ and $A$ as the tridiagonal matrix with diagonal $(1,\dots,1)$ and super- and sub-diagonals $(-\frac{1}{2},\dots,-\frac{1}{2})$. Next, we consider from [16], the case where $b=0$, AA is a diagonal matrix with entries $a_{ii}=\sin^{2}(\frac{\pi i}{2d})$, and $x_{0}=(1/a_{11},\dots,1/a_{dd})$. Finally, we consider the instance with $x_{0}=0$, $b=(-1,-2,\dots,-d)$ and $A$ as the diagonal matrix with diagonal $(1^{2},2^{2},\dots,d^{2})$. We refer to these problem instances as Problem A, Problem B, and Problem C, respectively. In Figure˜5, we show our methods remain competitive in these degenerate settings with $d=1000$.
@@ 5 Theory
$\epsilon(ce_{s},0)$ $=ca_{s}+\delta_{n}-\frac{L_{n}}{2}\|z_{s+1}-x_{0}\|^{2}=c\tau_{s}(f_{s}-\frac{1}{2L_{s}}\|g_{s}\|^{2}-v_{m})+\delta_{n}$
$=c\tau_{s}(v_{s}-v_{m})+\delta_{n}\geq 0$
@@
$=\kappa\tau_{s}f_{s}-\kappa\tau_{s}\frac{1}{2L_{s}}\|{g_{s}}\|^{2}-\frac{\kappa L_{s}}{2}\|{x_{0}}\|^{2}+\frac{\kappa L_{s}}{2}\|{z_{s+1}}\|^{2}-L_{n}c\langle{z_{s+1}-x_{0}},{x_{0}}\rangle$
$\qquad\qquad\qquad-\kappa v_{m}\tau_{s}+L_{n}\tau_{s}\left(\frac{1}{L_{s}^{2}}-\frac{1}{L_{n}^{2}}\right)\frac{1}{2}\|{g_{s}}\|^{2}-\frac{L_{n}}{2}\|{c(z_{s+1}-x_{0})}\|^{2}$
$\geq\kappa\tau_{s}f_{s}-\kappa\tau_{s}\frac{1}{2L_{n}}\|{g_{s}}\|^{2}-\kappa\tau_{s}v_{m}+\tau_{s}\left(\frac{1}{L_{s}}-\frac{1}{L_{n}}\right)\frac{1}{2}\|g_{s}\|^{2}$
$\qquad\qquad\qquad+\frac{L_{n}(c-c^{2})}{2}\|z_{s+1}-x_{0}\|^{2}$
@@
$U^{\prime}$ $=\left(\frac{L_{n}}{2}\|{x_{0}}\|^{2}-\frac{L_{n}}{2}\|{z^{\prime}}\|^{2}-\tau^{\prime}v_{m}+\Delta^{\prime}+\delta_{n}\right)+\tau^{\prime}f_{\star}+\langle L_{n}(z^{\prime}-x_{0}),x_{\star}\rangle,$
$U^{\prime}$ $=\left(-\sum_{i={n-k}}^{n-1}{\rho}_{i}{h}_{i}-\sum_{i={n-k}}^{n-1}{\gamma}_{i}{w}_{i}+\sum_{i={n-k}}^{n-1}{\rho}_{i}{\Delta}_{i}+\epsilon\right)+\left(\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}\right)f_{\star}$
$\qquad\qquad\qquad+\langle L_{n}Z\rho-L_{n}G\gamma,x_{\star}\rangle.$
@@
$\left\{\begin{array}[]{ll}\max_{\rho,\gamma}&\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}\\ \text{s.t.}&\frac{L_{n}}{2}\|{Z\rho-G\gamma}\|^{2}\leq\sum_{i=n-k}^{n-1}\rho_{i}(h_{i}-v_{m}\tau_{i})+\sum_{i=n-k}^{n-1}\gamma_{i}(w_{i}-v_{m})\\ &\qquad\qquad\qquad\qquad\qquad-\langle L_{n}Z\rho,x_{0}\rangle+\langle L_{n}G\gamma,x_{0}\rangle+\delta_{n}\\ &\rho,\,\gamma\geq 0.\end{array}\right$. (5.21)
Recalling our definition of $\epsilon(\rho,\gamma)$ from (3.11), this is exactly our optimization problem (3.12). Finally, with $\tau^{\prime}$ as the solution to (5.21), i.e. $\tau^{\prime}=\sum_{i={n-k}}^{n-1}{\rho}_{i}{\tau}_{i}+\sum_{i={n-k}}^{n-1}{\gamma}_{i}$, and $\tau_{n}$ set according to Algorithm˜3, we recover the exact form of Algorithm˜4.
@@
The final equality of our theorem $\frac{1}{\tau_{0,N}}=\frac{2}{N(N+1)+\sqrt{2N(N+1)}}$ comes directly from [46, Theorem 4].
@@
${\tau}^{(\ell)}_{n+1}\geq{\tau}^{(\ell)}_{n}$ $\geq\frac{2{L}^{(\ell)}_{n}}{\mu}+\frac{2{\Delta}^{(\ell)}_{n}}{f({x}^{(\ell)}_{0})-f({x}^{(\ell)}_{n})}$
$\geq\frac{2{L}^{(\ell)}_{n+1}}{\mu}+\frac{2{\Delta}^{(\ell)}_{n+1}}{f({x}^{(\ell)}_{0})-f(x_{\star})}$
$=\frac{2{L}^{(\ell)}_{n+1}(f({x}^{(\ell)}_{0})-f(x_{\star}))+2\mu{\Delta}^{(\ell)}_{n+1}}{\mu(f({x}^{(\ell)}_{0})-f(x_{\star}))}.$
@@
[…]
$\leq\left(\frac{{L}^{(\ell)}_{n+1}(f({x}^{(\ell)}_{0})-f(x_{\star}))+\mu{\Delta}^{(\ell)}_{n+1}}{\mu}\right)$